# Generative Deep Learning Framework for Inverse Design of Fuels

### v2

@@ 3 Regression Model Development for RON
Leveraging the latent space embeddings from the Co-VAE model, various regression algorithms were explored to determine the most effective model for RON prediction. This approach enables us to leverage the rich, informative features captured in the latent space while tailoring the regression model to best fit the specific characteristics of the RON prediction task. The study investigated a range of model classes—including ensemble-based methods, linear regression techniques, regularization, deep learning (TabNet) [22, 23], and SuperLearner approaches [11, 24, 25, 26]—each of which offers unique advantages for capturing diverse patterns and relationships in the data. Comprehensive hyperparameter optimization was performed for each model using NSGA-2 [27] from the Pymoo package [28]. This multi-objective evolutionary algorithm simultaneously optimized Coefficient of Determination (R²), MAE, and Root Mean Squared Error (RMSE), ensuring a balanced trade-off between predictive accuracy and generalization. A population size of 100 and 20 iterations were employed during the NSGA-2 optimization, which was carried out over the complete hyperparameter search space (see Table 3) with 10-fold cross-validation to mitigate overfitting. The candidate with the minimum MAE value was selected as the best candidate from the Pareto front. The candidate selection is mathematically formulated as follows:
Here, represents the MAE for the candidate, and denotes the optimal hyperparameter configuration corresponding to the minimum MAE value. Table 4 illustrates the performance of all evaluated models on the RON test set, where the CatBoost model emerged as the top performer for RON prediction based on latent space representations, achieving an R² of 0.929, an MAE of 5.365, and an RMSE of 8.090.
@@
To further ensure the robustness of the final model, an additional 10-fold cross-validation was performed on the entire RON dataset using the CatBoost model, configured with the best hyperparameters obtained from the multi-objective optimization. The final optimized model demonstrated a cross-validation performance of R² = 0.869 ± 0.102, MAE = 4.935 ± 1.041, and RMSE = 7.879 ± 2.964, as shown in the parity plot in Figure 4. This optimized model is the basis for the subsequent molecular screening described in the next section. By integrating the latent space embeddings from the Co-VAE with advanced regression techniques, the RON prediction accuracy was effectively improved compared to the initial property predictor within the Co-VAE. This two-step approach—first learning a rich latent representation through the Co-VAE and then fine-tuning a separate regression model—demonstrates the efficacy of combining deep generative models with traditional machine learning algorithms in the context of fuel property prediction. This methodology enhances the accuracy of property prediction and provides a flexible framework for exploring complex chemical spaces in fuel design. Decoupling the regression model allows for the inclusion of more sophisticated algorithms and optimization strategies, potentially leading to further improvements in predictive performance. Note that this does not necessarily mean the latent representation alone surpasses other feature formats or that CatBoost is universally superior to neural networks; rather, it illustrates how the Co-VAE’s generative capabilities, when paired with a diverse set of predictive models, can effectively facilitate targeted fuel design.
@@ 4 Generation of Promising High-RON Fuel Molecules
To identify latent vectors corresponding to molecules with predicted RON values above 110, the Differential Evolution (DE) algorithm—a population-based, stochastic optimization method designed for continuous and high-dimensional spaces —is employed. DE utilizes mutation, crossover, and selection, inspired by natural evolution, to maximize the predicted RON from the regression model, subject to defined bounds of the latent space. The SciPy implementation of DE is leveraged to peform this optimization [29]. Once high-predicted RON latent vectors were identified, they were decoded into molecular structures using the decoder component of the Co-VAE model, which transforms each latent vector into a sequence of tokens representing a SMILES string. The generated SMILES strings underwent a two-step validation process to ensure chemical feasibility and consistency in the predicted performance. First, the RDKit, an open-source cheminformatics library [30], was used to verify that each SMILES string adheres to chemical valency rules and proper syntax, discarding any that fail this test. Beyond this initial screening, each decoded molecule was re-encoded using the Co-VAE encoder and re-evaluated with the regression model to confirm if its predicted RON remains above 110. This extra validation is essential because the VAE generates latent representations by sampling from a distribution, meaning the latent vector passed to the decoder is effectively a perturbed version of its ideal representation. Such perturbations can lead to fluctuations in the predicted RON, even if the molecule is chemically sound. This inherent variability—central to the variational approach—broadens chemical space exploration and necessitates a dual screening process to ensure robust, high-performance fuel candidates. This robust validation framework filters out decoding artifacts, ensuring that only reliable, high-performance fuel candidates progress to further experimental evaluation. The complete workflow for the fuel molecule generation is shown in Figure 5.
@@ 5 Conclusions
To enhance and expand the proposed methodology, future work will focus on fine-tuning the downselection process of the GDB-13 database to curate a dataset enriched with fuel-like species, potentially improving the relevance and applicability of the generated molecules. Leveraging transfer learning techniques, the Co-VAE can be initially trained on an extensive database with abundant property data before being fine-tuned for the RON prediction task, benefiting from the broader chemical knowledge while specializing in fuel properties. Implementing multi-task learning approaches, which simultaneously predict multiple properties within the regression model, can capture interrelated property trends and enhance overall predictive performance. Furthermore, by embedding UQ methods [8], the model could not only flag candidates with high predictive confidence—ideal for selection and experimental validation—but also identify regions of the latent space with higher uncertainty. These uncertain compounds can guide further data collection and model refinement, ensuring that exploration focuses on promising and underrepresented chemical spaces. Such an approach would refine the screening process, evolving the framework into a more robust and versatile tool for advanced fuel design, thereby contributing to the development of cleaner and more efficient energy sources. In addition, it is noted that standard methods for measuring RON, have experimental repeatability ranges of approximately ±1 RON. The data-driven approach currently exceeds this threshold, reinforcing the need for larger and more diverse datasets to reduce predictive uncertainty. Lastly, extending this framework to multi-component blends constitutes an important next step, as real engines commonly operate with complex fuel mixtures exhibiting non-linear synergistic effects.
Acknowledgments

### v3

@@ 3 Regression Model Development for RON
Leveraging the latent space embeddings from the Co-VAE model, various regression algorithms were explored to determine the most effective model for RON prediction. This approach enables us to leverage the rich, informative features captured in the latent space while tailoring the regression model to best fit the specific characteristics of the RON prediction task. The study investigated a range of model classes—including ensemble-based methods, linear regression techniques, regularization, deep learning (TabNet) 24, 25, and SuperLearner approaches 12, 26, 27, 28—each of which offers unique advantages for capturing diverse patterns and relationships in the data. Comprehensive hyperparameter optimization was performed for each model using NSGA-2 29 from the Pymoo package 30. This multi-objective evolutionary algorithm simultaneously optimized Coefficient of Determination (R²), MAE, and Root Mean Squared Error (RMSE), ensuring a balanced trade-off between predictive accuracy and generalization. A population size of 100 and 20 iterations were employed during the NSGA-2 optimization, which was carried out over the complete hyperparameter search space (see Table 3) with 10-fold cross-validation to mitigate overfitting. The candidate with the minimum MAE value was selected as the best candidate from the Pareto front. The candidate selection is mathematically formulated as follows:
Here, represents the MAE for the candidate, and denotes the optimal hyperparameter configuration corresponding to the minimum MAE value. Table 4 illustrates the performance of all evaluated models on the RON test set, where the CatBoost model emerged as the top performer for RON prediction based on latent space representations, achieving an R² of 0.929, an MAE of 5.365, and an RMSE of 8.090. This performance is on par with recent literature reporting MAE values of 5.22 7 and 4.86 13, though Chen et al. 13 work achieved 3.71 MAE by incorporating additional branching ratio feature, which could potentially be integrated into our framework for future improvements.
@@
To further ensure the robustness of the final model, an additional 10-fold cross-validation was performed on the entire RON dataset using the CatBoost model, configured with the best hyperparameters obtained from the multi-objective optimization. The final optimized model demonstrated a cross-validation performance of R² = 0.869 ± 0.102, MAE = 4.935 ± 1.041, and RMSE = 7.879 ± 2.964, as shown in the parity plot in Figure 4. The worst-performing outlier is n-propyl cyclohexane and exhibits a predicted RON value significantly higher than the experimental value. This optimized model is the basis for the subsequent molecular screening described in the next section. By integrating the latent space embeddings from the Co-VAE with advanced regression techniques, the RON prediction accuracy was effectively improved compared to the initial property predictor within the Co-VAE. This two-step approach—first learning a rich latent representation through the Co-VAE and then fine-tuning a separate regression model—demonstrates the efficacy of combining deep generative models with traditional machine learning algorithms in the context of fuel property prediction. This methodology enhances the accuracy of property prediction and provides a flexible framework for exploring complex chemical spaces in fuel design. Decoupling the regression model allows for the inclusion of more sophisticated algorithms and optimization strategies, potentially leading to further improvements in predictive performance. Note that this does not necessarily mean the latent representation alone surpasses other feature formats or that CatBoost is universally superior to neural networks; rather, it illustrates how the Co-VAE’s generative capabilities, when paired with a diverse set of predictive models, can effectively facilitate targeted fuel design.
@@ 4 Generation of Promising High-RON Fuel Molecules
To identify latent vectors corresponding to molecules with predicted RON values above 110, the Differential Evolution (DE) algorithm—a population-based, stochastic optimization method designed for continuous and high-dimensional spaces —is employed. DE utilizes mutation, crossover, and selection, inspired by natural evolution, to maximize the predicted RON from the regression model, subject to defined bounds of the latent space. The SciPy implementation of DE is leveraged to peform this optimization 31. Once high-predicted RON latent vectors were identified, they were decoded into molecular structures using the decoder component of the Co-VAE model, which transforms each latent vector into a sequence of tokens representing a SMILES string. The generated SMILES strings underwent a two-step validation process to ensure chemical feasibility and consistency in the predicted performance. First, the RDKit, an open-source cheminformatics library 32, was used to verify that each SMILES string adheres to chemical valency rules and proper syntax, discarding any that fail this test. Beyond this initial screening, each decoded molecule was re-encoded using the Co-VAE encoder and re-evaluated with the regression model to confirm if its predicted RON remains above 110. This extra validation is essential because the VAE generates latent representations by sampling from a distribution, meaning the latent vector passed to the decoder is effectively a perturbed version of its ideal representation. Such perturbations can lead to fluctuations in the predicted RON, even if the molecule is chemically sound. This inherent variability—central to the variational approach—broadens chemical space exploration and necessitates a dual screening process to ensure robust, high-performance fuel candidates. This robust validation framework filters out decoding artifacts, ensuring that only reliable, high-performance fuel candidates progress to further experimental evaluation. The complete workflow for the fuel molecule generation is shown in Figure 5.
@@ 5 Conclusions
To enhance and expand the proposed methodology, future work will focus on fine-tuning the downselection process of the GDB-13 database to curate a dataset enriched with fuel-like species, potentially improving the relevance and applicability of the generated molecules. Leveraging transfer learning techniques, the Co-VAE can be initially trained on an extensive database with abundant property data before being fine-tuned for the RON prediction task, benefiting from the broader chemical knowledge while specializing in fuel properties. Implementing multi-task learning approaches, which simultaneously predict multiple properties within the regression model, can capture interrelated property trends and enhance overall predictive performance. Furthermore, by embedding UQ methods 9, the model could not only flag candidates with high predictive confidence—ideal for selection and experimental validation—but also identify regions of the latent space with higher uncertainty. These uncertain compounds can guide further data collection and model refinement, ensuring that exploration focuses on promising and underrepresented chemical spaces. Such an approach would refine the screening process, evolving the framework into a more robust and versatile tool for advanced fuel design, thereby contributing to the development of cleaner and more efficient energy sources. In addition, it is noted that standard methods for measuring RON, have experimental repeatability ranges of approximately ±1 RON. The data-driven approach currently exceeds this threshold, reinforcing the need for larger and more diverse datasets to reduce predictive uncertainty. Lastly, extending this framework to multi-component blends constitutes an important next step, as real engines commonly operate with complex fuel mixtures exhibiting non-linear synergistic effects.